rawrxd:

What is the slope of the equation mc016-1.jpg?

umm:

Welcome to QuestionCove! Is there an image included for the slope?

Angle:

You can use the "Attach File" button below to try to post your mc016-1.jpg picture we are missing some information

rawrxd:

Angle:

awesome :) so here, we can see the equation in what is called "slope intercept form" where \(y = mx + b\) m =slope what number is sitting where the "m" is?

Angle:

\(y = \large \frac{5}{4}x - \frac{7}{4}\) ^